OpenStudy (anonymous):

To get to school each day, Jamie travels 3 miles south, turns left and goes 4 miles east, arriving at the school door. What is the straight line distance between his house and the school?

OpenStudy (denonakavro):

3^2 + 4^2 = c^2 9+16 =25 sqrt25 =5